Show that the series 8/(n⁴+x⁴) is uniformly convergent for all real values of x. n=1 to infinity .

1
Expert's answer
2021-12-08T10:27:59-0500

since "|\frac{8}{n^4+x^4}|\le \frac{8}{n^4}" and series "\sum \frac{8}{n^4}" converges, then series "\sum \frac{8}{n^4+x^4}" converges as well